### v1

@@ 4.4 Coordinate ascent variational inference
In order to maximize the ELBO $\mathcal{J}(q(\mathcal{S},\mathcal{L},\boldsymbol{\theta}))$, we apply coordinate ascent variational inference [7, 45]. Starting from an initial density $(q_{0}(\boldsymbol{\theta}),q_{0}(\mathcal{L}),q_{0}(\mathcal{S}))$, the densities of $\mathcal{X}$, $\mathcal{S}$ and $\theta$ are updated as follows:
(26) $q_{\ell}(\mathcal{S})$ $=$ $\underset{q(\mathcal{S})\in\mathcal{Q}_{\mathcal{N}}}{\operatorname{argmax}}\mathcal{J}\left(q(\mathcal{S})q_{\ell-1}(\mathcal{L})q_{\ell-1}(\boldsymbol{\theta})\right),$
(27) $q_{\ell}(\mathcal{L})$ $=$ $\underset{q(\mathcal{L})\in\mathcal{Q}_{\mathcal{N}}}{\operatorname{argmax}}\mathcal{J}\left(q_{\ell}(\mathcal{S})q(\mathcal{L})q_{\ell-1}(\boldsymbol{\theta})\right),$
(28) $q_{\ell}(\boldsymbol{\theta})$ $=$ $\underset{q(\boldsymbol{\theta})\in\mathcal{Q}_{\mathcal{G}}}{\operatorname{argmax}}\mathcal{J}\left(q_{\ell}(\mathcal{L})q_{\ell}(\mathcal{S})q(\boldsymbol{\theta})\right),$
@@ 4.4.1 The density $q_{\ell}(\mathcal{S})$
[…]
@@
$=$ $-\sum_{ijk}\left(\frac{\mathsf{E}_{\theta_{1}}^{\ell-1}}{2}\left(\mathcal{X}_{ijk}-\mathsf{E}_{\mathcal{L}_{ijk}}^{\ell-1}-\mathcal{S}_{ijk}\right)^{2}+\mathsf{E}_{\theta_{2}}^{\ell-1}\left|\mathcal{S}_{ijk}\right|\right)+\text{const}.$
According to the discussion in Section 4.3.1, the mean of $\mathcal{S}_{ijk}$ is given by
@@
(30) $\mathsf{E}_{\mathcal{S}_{ijk}}^{\ell}=\left\{\begin{array}[]{ll}\mathcal{X}_{ijk}-\mathsf{E}_{\mathcal{L}_{ijk}}^{\ell-1}-\frac{\mathsf{E}_{\theta_{2}}^{\ell-1}}{\mathsf{E}_{\theta_{1}}^{\ell-1}},&\mbox{if }\mathcal{X}_{ijk}-\mathsf{E}_{\mathcal{L}_{ijk}}^{\ell-1}\geq\frac{\mathsf{E}_{\theta_{2}}^{\ell-1}}{\mathsf{E}_{\theta_{1}}^{\ell-1}},\\ \mathcal{X}_{ijk}-\mathsf{E}_{\mathcal{L}_{ijk}}^{\ell-1}+\frac{\mathsf{E}_{\theta_{2}}^{\ell-1}}{\mathsf{E}_{\theta_{1}}^{\ell-1}},&\mbox{if }\mathcal{X}_{ijk}-\mathsf{E}_{\mathcal{L}_{ijk}}^{\ell-1}\leq-\frac{\mathsf{E}_{\theta_{2}}^{\ell-1}}{\mathsf{E}_{\theta_{1}}^{\ell-1}},\\ 0,&\mbox{others}.\\ \end{array}\right.$
@@
(31) $q(\mathcal{S}_{ijk})=\mathcal{N}(\mathcal{S}|\mathsf{E}_{\mathcal{S}_{ijk}}^{\ell},\boldsymbol{\Sigma}_{\mathcal{S}_{ijk}}^{\ell}).$
@@ 4.4.2 The density $q_{\ell}(\mathcal{L})$
[…]
@@
$\mathbb{E}_{q_{\ell}(\mathcal{S})q_{\ell-1}(\boldsymbol{\theta})}\left[\log p(\mathcal{X},\mathcal{S},\mathcal{L},\boldsymbol{\theta})\right]=-\left(\tfrac{\mathsf{E}_{\theta_{1}}^{\ell-1}}{2}\left\|\mathcal{L}-(\mathcal{X}-\mathsf{E}_{\mathcal{S}}^{\ell})\right\|_{F}^{2}+\mathsf{E}_{\theta_{3}}^{\ell-1}\left\|\mathcal{L}\right\|_{*}\right)+\text{const}.$
@@
(33) $\mathsf{E}_{\mathcal{L}}^{\ell}=\mathcal{U}^{\ell}*\mathcal{D}^{\ell}_{\tau}*{\mathcal{V}^{\ell}}^{T},$
@@
We apply (22) and then obtain the covariance matrix of the vector ¯$\overline{\mathcal{L}}_{:jk}$
$\boldsymbol{\Sigma}_{\overline{\mathcal{L}}_{:jk}}^{\ell}=\overline{\mathcal{U}}_{::k}^{\ell}{\overline{\mathcal{D}}_{\tau}^{\ell}}_{::k}\left(\mathsf{E}_{\theta_{1}}^{\ell-1}{\overline{\mathcal{D}}_{\tau}^{\ell}}_{::k}+{\mathsf{E}_{\theta_{3}}^{\ell-1}}\mathbf{I}\right)^{-1}{\overline{\mathcal{U}}_{::k}^{\ell}}^{T}.$
@@
(34) $q(\overline{\mathcal{L}})=\prod_{jk}\mathcal{N}\left(\overline{\mathcal{L}}_{:jk}|\mathsf{E}_{\overline{\mathcal{L}}_{:jk}}^{\ell},\boldsymbol{\Sigma}_{\overline{\mathcal{L}}_{:jk}}^{\ell}\right).$
@@ 4.4.3 The density $q_{\ell}(\boldsymbol{\theta})$
[…]
where $\mathcal{Q}_{\mathcal{G}}$ is the set of all the density functions for Gamma distribution. By taking the partial derivative of the objective function in (35) with respect to $q(\boldsymbol{\theta})$, and letting it be equal to 0, we obtain the optimal $q(\boldsymbol{\theta})$ is proportional to
@@
To facilitate the interpretation of the $\theta_{i}$ parameters within a Bayesian framework, we consider assigning Gamma priors to these parameters. By comparing the coefficients in $\mathbb{E}_{q_{\ell}(\mathcal{S},\mathcal{L})}\log p(\mathcal{X},\mathcal{S},\mathcal{L},\boldsymbol{\theta})$ with those of a Gamma density ($\mathcal{G}(x|a,b)\propto x^{a-1}\exp(-bx)$), we can infer the shape $a_{\theta_{i}}^{\ell}$ and scale $b_{\theta_{i}}^{\ell}$ parameters for each $\theta_{i}$. Consequently, the shape parameters are given by:
@@
(36) $\left\{\begin{array}[]{l}b_{\theta_{1}}^{\ell}=\frac{1}{2}\mathbb{E}_{q_{\ell}(\mathcal{L})q_{\ell}(\mathcal{S})}\left[\left\|\mathcal{X}-\mathcal{S}-\mathcal{L}\right\|_{F}^{2}\right],\\ b_{\theta_{2}}^{\ell}=\mathbb{E}_{q_{\ell}(\mathcal{S})}\left\|\mathcal{S}\right\|_{1},\\ b_{\theta_{3}}^{\ell}=\mathbb{E}_{q_{\ell}(\mathcal{L})}\left\|\mathcal{L}\right\|_{*}.\end{array}\right.$
@@
For the expectation of the nuclear norm $\left\|\mathcal{L}\right\|_{*}$, it is the arithmetic mean of each slice ¯$\overline{\mathcal{L}}_{::k}$ of the tensor $\mathcal{L}$. Hence we need to evaluate $\mathbb{E}_{q_{\ell}(\overline{\mathcal{L}}_{::k})}\|\overline{\mathcal{L}}_{::k}\|_{*}$. According to (25), we have
@@
Now, we focus on the expectation of the nuclear norm $\left\|\mathcal{L}\right\|_{*}$, which requires evaluating $\mathbb{E}_{q_{\ell}(\overline{\mathcal{L}}_{::k})}\|\overline{\mathcal{L}}_{::k}\|_{*}$ for each slice ¯$\overline{\mathcal{L}}_{::k}$ of the tensor $\mathcal{L}$.
Finally, we summarize the proposed adaptive method in Algorithm 1. For simplicity, we refer to our proposed algorithm for solving the tensor nuclear norm model as VBITNN.
@@
3:  while $\ell\leq\ell\rm{Max}$ or not converged do
4:     $\mathsf{E}_{\mathcal{S}_{ijk}}^{\ell}=\left\{\begin{array}[]{ll}\mathcal{X}_{ijk}-\mathsf{E}_{\mathcal{L}_{ijk}}^{\ell-1}-\frac{\mathsf{E}_{\theta_{2}}^{\ell-1}}{\mathsf{E}_{\theta_{1}}^{\ell-1}},&\mbox{if }\mathcal{X}_{ijk}-\mathsf{E}_{\mathcal{L}_{ijk}}^{\ell-1}\geq\frac{\mathsf{E}_{\theta_{2}}^{\ell-1}}{\mathsf{E}_{\theta_{1}}^{\ell-1}}\\ \mathcal{X}_{ijk}-\mathsf{E}_{\mathcal{L}_{ijk}}^{\ell-1}+\frac{\mathsf{E}_{\theta_{2}}^{\ell-1}}{\mathsf{E}_{\theta_{1}}^{\ell-1}},&\mbox{if }\mathcal{X}_{ijk}-\mathsf{E}_{\mathcal{L}_{ijk}}^{\ell-1}\leq-\frac{\mathsf{E}_{\theta_{2}}^{\ell-1}}{\mathsf{E}_{\theta_{1}}^{\ell-1}}\\ 0,&\mbox{others }\\ \end{array}\right.$
@@
7:     $\boldsymbol{\Sigma}_{\mathcal{S}_{ijk}}^{\ell}=\frac{\mathsf{E}_{\theta_{1}}^{\ell-1}|\mathsf{E}_{\mathcal{S}_{ijk}}^{\ell}|}{\mathsf{E}_{\theta_{1}}^{\ell-1}|\mathsf{E}_{\mathcal{S}_{ijk}}^{\ell}|+\mathsf{E}_{\theta_{2}}^{\ell-1}},$ and $\boldsymbol{\Sigma}_{\overline{\mathcal{L}}_{:jk}}^{\ell}=\overline{\mathcal{U}}_{::k}^{\ell}{\overline{\mathcal{D}}_{\tau}^{\ell}}_{::k}\left(\mathsf{E}_{\theta_{1}}^{\ell-1}{\overline{\mathcal{D}}_{\tau}^{\ell}}_{::k}+{\mathsf{E}_{\theta_{3}}^{\ell-1}}\mathbf{I}\right)^{-1}{\overline{\mathcal{U}}_{::k}^{\ell}}^{T}$
@@ 4.5 Variational Bayesian inference for weighted tensor nuclear norm
In this subsection, we consider a variant of the tensor nuclear norm by reweighting the singular values [22, 11]. Note that the standard tensor nuclear norm can be regarded as the special version of the weighted tensor nuclear norm, where the weighting matrix consists of elements that are all equal to one. Formally, for a non-negative matrix $\mathbf{W}\in\mathbb{R}^{\min(n_{1},n_{2})\times n_{3}}$ with column vectors $\mathbf{W}_{:k}$, the weighted tensor nuclear norm $\left\|\mathcal{A}\right\|_{\mathbf{W}*}$ is defined as:
@@
(37) $\min_{\mathcal{S},\mathcal{L}}\left\{\frac{\theta_{1}}{2}\|\mathcal{X}-\mathcal{L}-\mathcal{S}\|_{F}^{2}+\theta_{2}\|\mathcal{S}\|_{1}+\theta_{3}\|\mathcal{L}\|_{\mathbf{W}*}\right\}.$
During the inference of $\mathcal{L}$, we update the expectation of ¯$\overline{\mathcal{L}}_{::k}$ in (34) to:
(38) $\mathsf{E}_{\mathcal{L}}^{\ell}=\mathcal{U}^{\ell}*\mathcal{D}^{\ell}_{\mathbf{W}}*{\mathcal{V}^{\ell}}^{T},$
@@
Concurrently, the covariance matrix of ¯$\overline{\mathcal{L}}_{::k}$ is adjusted to:
$\boldsymbol{\Sigma}_{\overline{\mathcal{L}}_{:jk}}^{\ell}=\overline{\mathcal{U}}_{::k}^{\ell}{\overline{\mathcal{D}}^{\ell}}_{::k}(\tau^{\ell})\left(\mathsf{E}_{\theta_{1}}^{\ell-1}{\overline{\mathcal{D}}_{\tau}^{\ell}}_{::k}+\mathsf{E}_{\theta_{3}}^{\ell-1}\operatorname{diag}(\mathbf{W}_{:k})\right)^{-1}{\overline{\mathcal{U}}_{::k}^{\ell}}^{T}.$

### v2

@@ 4.4 Coordinate ascent variational inference
In order to maximize the ELBO $\mathcal{J}(q(\mathcal{S},\mathcal{L},\boldsymbol{\theta}))$, we apply coordinate ascent variational inference (CAVI) [8, 47]. Starting from an initial density $(q_{0}(\boldsymbol{\theta}),q_{0}(\mathcal{L}),q_{0}(\mathcal{S}))$, the densities of $\mathcal{S}$, $\mathcal{L}$ and $\boldsymbol{\theta}$ are updated as follows:
$q_{\ell}(\mathcal{S})$ $=$ $\underset{q(\mathcal{S})\in\mathcal{Q}_{\mathcal{N}}}{\operatorname{argmax}}\mathcal{J}\left(q(\mathcal{S})q_{\ell-1}(\mathcal{L})q_{\ell-1}(\boldsymbol{\theta})\right),$ (26)
$q_{\ell}(\mathcal{L})$ $=$ $\underset{q(\mathcal{L})\in\mathcal{Q}_{\mathcal{N}}}{\operatorname{argmax}}\mathcal{J}\left(q_{\ell}(\mathcal{S})q(\mathcal{L})q_{\ell-1}(\boldsymbol{\theta})\right),$ (27)
$q_{\ell}(\boldsymbol{\theta})$ $=$ $\underset{q(\boldsymbol{\theta})\in\mathcal{Q}_{\mathcal{G}}}{\operatorname{argmax}}\mathcal{J}\left(q_{\ell}(\mathcal{S})q_{\ell}(\mathcal{L})q(\boldsymbol{\theta})\right),$ (28)
@@ 4.4.1 The density $q_{\ell}(\mathcal{S})$
$\mathop{\rm argmax}_{q(\mathcal{S})\in\mathcal{Q}_{\mathcal{N}}}\mathcal{J}(q(\mathcal{S}),q_{\ell-1}(\mathcal{L},\boldsymbol{\theta}))=\mathop{\rm argmax}_{q(\mathcal{S})\in\mathcal{Q}_{\mathcal{N}}}\int q(\mathcal{S}){\mathbb{E}_{q_{\ell-1}(\mathcal{L},\boldsymbol{\theta})}\log\frac{p(\mathcal{X},\mathcal{S},\mathcal{L},\boldsymbol{\theta})}{q(\mathcal{S})}}d\mathcal{S}.$ (29)
@@
$=$ $-\sum_{ijk}\left(\frac{\mathsf{E}_{\theta_{1}}^{\ell-1}}{2}\left(\mathcal{X}_{ijk}-\mathsf{E}_{\mathcal{L}_{ijk}}^{\ell-1}-\mathcal{S}_{ijk}\right)^{2}+\mathsf{E}_{\theta_{2}}^{\ell-1}\left|\mathcal{S}_{ijk}\right|\right)+\text{const},$
where const is a constant independent of $\mathcal{S}$. According to the discussion in Section 4.3.1, the mean of $\mathcal{S}_{ijk}$ is given by
@@
[…]
@@
$q_{\ell}(\mathcal{S}_{ijk})=\mathcal{N}(\mathcal{S}|\mathsf{E}_{\mathcal{S}_{ijk}}^{\ell},\boldsymbol{\Sigma}_{\mathcal{S}_{ijk}}^{\ell}).$ (31)
@@ 4.4.2 The density $q_{\ell}(\mathcal{L})$
$\mathop{\rm argmax}_{q(\mathcal{L})\in\mathcal{Q}_{\mathcal{N}}}\mathcal{J}(q_{\ell}(\mathcal{S}),q(\mathcal{L}),q_{\ell-1}(\boldsymbol{\theta}))=\mathop{\rm argmax}_{q(\mathcal{L})\in\mathcal{Q}_{\mathcal{N}}}\int q(\mathcal{L}){\mathbb{E}_{q_{\ell}(\mathcal{S})q_{\ell-1}(\boldsymbol{\theta})}\log\tfrac{p(\mathcal{X},\mathcal{S},\mathcal{L},\boldsymbol{\theta})}{q(\mathcal{L})}}d\mathcal{L}.$ (32)
@@
$\begin{split}\mathbb{E}_{q_{\ell}(\mathcal{S})q_{\ell-1}(\boldsymbol{\theta})}\left[\log p(\mathcal{X},\mathcal{S},\mathcal{L},\boldsymbol{\theta})\right]&=-\left(\frac{\mathsf{E}_{\theta_{1}}^{\ell-1}}{2}\left\|\mathcal{L}-(\mathcal{X}-\mathsf{E}_{\mathcal{S}}^{\ell})\right\|_{F}^{2}+\mathsf{E}_{\theta_{3}}^{\ell-1}\left\|\mathcal{L}\right\|_{*}\right)+\text{const}\\ &=-\left(\frac{\mathsf{E}_{\theta_{1}}^{\ell-1}}{2n_{3}}\left\|\overline{\mathcal{L}}-(\overline{\mathcal{X}}-\overline{\mathsf{E}_{\mathcal{S}}^{\ell}})\right\|_{F}^{2}+\frac{\mathsf{E}_{\theta_{3}}^{\ell-1}}{n_{3}}\left\|\overline{\mathcal{L}}\right\|_{*}\right)+\text{const}\end{split}$
@@
$\mathsf{E}_{\mathcal{L}}^{\ell}=\mathcal{U}^{\ell}*\mathcal{D}^{\ell}_{\tau}*{\mathcal{V}^{\ell}}^{T},$ (33)
@@
We apply (21) and then obtain the covariance matrix of the vector $\overline{\mathcal{L}}_{:jk}$
$\boldsymbol{\Sigma}_{\overline{\mathcal{L}}_{:jk}}^{\ell}=n_{3}\overline{\mathcal{U}}_{::k}^{\ell}{\overline{\mathcal{D}}_{\tau}^{\ell}}_{::k}\left(\mathsf{E}_{\theta_{1}}^{\ell-1}{\overline{\mathcal{D}}_{\tau}^{\ell}}_{::k}+{\mathsf{E}_{\theta_{3}}^{\ell-1}}\mathbf{I}\right)^{-1}{\overline{\mathcal{U}}_{::k}^{\ell}}^{T}.$
@@
$q_{\ell}(\overline{\mathcal{L}})=\prod_{jk}\mathcal{N}\left(\overline{\mathcal{L}}_{:jk}|\mathsf{E}_{\overline{\mathcal{L}}_{:jk}}^{\ell},\boldsymbol{\Sigma}_{\overline{\mathcal{L}}_{:jk}}^{\ell}\right).$ (34)
@@ 4.4.3 The density $q_{\ell}(\boldsymbol{\theta})$
$\mathop{\rm argmax}_{q(\boldsymbol{\theta})\in\mathcal{Q}_{\mathcal{G}}}\mathcal{J}(q_{\ell}(\mathcal{S}),q_{\ell}(\mathcal{L}),q(\boldsymbol{\theta}))=\mathop{\rm argmax}_{q(\boldsymbol{\theta})\in\mathcal{Q}_{\mathcal{G}}}\int q(\boldsymbol{\theta})\mathbb{E}_{q_{\ell}(\mathcal{S},\mathcal{L})}\log\tfrac{p(\mathcal{X},\mathcal{S},\mathcal{L},\boldsymbol{\theta})}{q(\boldsymbol{\theta})}d\boldsymbol{\theta},$ (35)
where $\mathcal{Q}_{\mathcal{G}}$ is the set of all the density functions for the Gamma distribution. By taking the partial derivative of the objective function in (35) with respect to $q(\boldsymbol{\theta})$, and letting it be equal to 0, we obtain that the optimal $q(\boldsymbol{\theta})$ is proportional to
@@
Since $q(\boldsymbol{\theta})=\prod_{i=1}^{3}q(\theta_{i})$ is assumed to factorize due to the independence of the model components, we derive the form of each $q(\theta_{i})$ by comparing the coefficients of $\theta_{i}$ and $\log\theta_{i}$ with the log-density of a Gamma distribution
$\log q(\boldsymbol{\theta})=\sum_{i=1}^{3}\left((a_{\theta_{i}}^{\ell}-1)\log\theta_{i}-b_{\theta_{i}}^{\ell}\theta_{i}\right)+\text{const},$
where $a_{\theta_{i}}^{\ell}$ and $b_{\theta_{i}}^{\ell}$ are the shape and rate parameters, respectively. By comparing the coefficients in $\mathbb{E}_{q_{\ell}(\mathcal{S},\mathcal{L})}\log p(\mathcal{X},\mathcal{S},\mathcal{L},\boldsymbol{\theta})$ with those of a Gamma density ($\mathcal{G}(x|a,b)\propto x^{a-1}\exp(-bx)$), we can infer the shape $a_{\theta_{i}}^{\ell}$ and scale $b_{\theta_{i}}^{\ell}$ parameters for each $\theta_{i}$. Consequently, the shape parameters are given by:
@@
$\left\{\begin{array}[]{l}b_{\theta_{1}}^{\ell}=\frac{1}{2}\mathbb{E}_{q_{\ell}(\mathcal{L})q_{\ell}(\mathcal{S})}\left[\left\|\mathcal{X}-\mathcal{S}-\mathcal{L}\right\|_{F}^{2}\right],\\ b_{\theta_{2}}^{\ell}=\mathbb{E}_{q_{\ell}(\mathcal{S})}\left\|\mathcal{S}\right\|_{1},\\ b_{\theta_{3}}^{\ell}=\mathbb{E}_{q_{\ell}(\mathcal{L})}\left\|\mathcal{L}\right\|_{*}.\end{array}\right$. (36)
@@
For the expectation of the nuclear norm $\left\|\mathcal{L}\right\|_{*}$, it is the arithmetic mean of each slice $\overline{\mathcal{L}}_{::k}$ of the tensor $\mathcal{L}$. Hence we need to evaluate $\mathbb{E}_{q_{\ell}(\overline{\mathcal{L}}_{::k})}\|\overline{\mathcal{L}}_{::k}\|_{*}$. According to (25), we have
@@
Now, we focus on the expectation of the nuclear norm $\left\|\mathcal{L}\right\|_{*}$, which requires evaluating $\mathbb{E}_{q_{\ell}(\overline{\mathcal{L}}_{::k})}\|\overline{\mathcal{L}}_{::k}\|_{*}$ for each slice $\overline{\mathcal{L}}_{::k}$ of the tensor $\mathcal{L}$.
We summarize the proposed adaptive method in Algorithm 1. For simplicity, we refer to our proposed algorithm for solving the tensor nuclear norm model as VBITNN.
@@
3:while $\ell\leq\ell_{\rm{Max}}$ or not converged do
[…]
@@
7:  $\boldsymbol{\Sigma}_{\mathcal{S}_{ijk}}^{\ell}=\frac{\mathsf{E}_{\theta_{1}}^{\ell-1}|\mathsf{E}_{\mathcal{S}_{ijk}}^{\ell}|}{\mathsf{E}_{\theta_{1}}^{\ell-1}|\mathsf{E}_{\mathcal{S}_{ijk}}^{\ell}|+\mathsf{E}_{\theta_{2}}^{\ell-1}},$ and $\boldsymbol{\Sigma}_{\overline{\mathcal{L}}_{:jk}}^{\ell}=n_{3}\overline{\mathcal{U}}_{::k}^{\ell}{\overline{\mathcal{D}}_{\tau}^{\ell}}_{::k}\left(\mathsf{E}_{\theta_{1}}^{\ell-1}{\overline{\mathcal{D}}_{\tau}^{\ell}}_{::k}+{\mathsf{E}_{\theta_{3}}^{\ell-1}}\mathbf{I}\right)^{-1}{\overline{\mathcal{U}}_{::k}^{\ell}}^{T}$
@@
Remark 4.1.
According to [7], a general theoretical treatment of analyzing the convergence of CAVI is missing in the literature. This is due to the lack of tractability of the updating formula involving unwieldy normalization constants and the technical challenge of dealing with optimization over infinite-dimensional distributions. Here, we will empirically show the convergence in Section 5.
@@ 4.5 Variational Bayesian inference for weighted tensor nuclear norm
In this subsection, we consider a variant of the tensor nuclear norm by reweighting the singular values [23, 12]. Note that the standard tensor nuclear norm can be regarded as a special version of the weighted tensor nuclear norm, where the weighting matrix consists of elements that are all equal to one. Formally, for a non-negative matrix $\mathbf{W}\in\mathbb{R}^{\min(n_{1},n_{2})\times n_{3}}$ with column vectors $\mathbf{W}_{:k}$, the weighted tensor nuclear norm $\left\|\mathcal{A}\right\|_{\mathbf{W}*}$ is defined as:
@@
$\min_{\mathcal{S},\mathcal{L}}\left\{\frac{\theta_{1}}{2}\|\mathcal{X}-\mathcal{L}-\mathcal{S}\|_{F}^{2}+\theta_{2}\|\mathcal{S}\|_{1}+\theta_{3}\|\mathcal{L}\|_{\mathbf{W}*}\right\}.$ (37)
During the inference of $\mathcal{L}$, we update the expectation of $\overline{\mathcal{L}}_{::k}$ in (34) to:
$\mathsf{E}_{\mathcal{L}}^{\ell}=\mathcal{U}^{\ell}*\mathcal{D}^{\ell}_{\mathbf{W}}*{\mathcal{V}^{\ell}}^{T},$ (38)
@@
Concurrently, the covariance matrix of $\overline{\mathcal{L}}_{::k}$ is adjusted to:
$\boldsymbol{\Sigma}_{\overline{\mathcal{L}}_{:jk}}^{\ell}=n_{3}\overline{\mathcal{U}}_{::k}^{\ell}{\overline{\mathcal{D}}^{\ell}}_{::k}(\tau^{\ell})\left(\mathsf{E}_{\theta_{1}}^{\ell-1}{\overline{\mathcal{D}}_{\tau}^{\ell}}_{::k}+\mathsf{E}_{\theta_{3}}^{\ell-1}\operatorname{diag}(\mathbf{W}_{:k})\right)^{-1}{\overline{\mathcal{U}}_{::k}^{\ell}}^{T}.$